Find the absolute extrema of the function on the indicated

interval


1. 𝑔(𝑑) = 2 sec 1/2 𝑑 ; [ βˆ’1/3πœ‹,1/2πœ‹ ]


2. 𝑔(π‘₯) = π‘₯Β³ + 3π‘₯Β² βˆ’ 9π‘₯ ; [βˆ’4, 4]


3. 𝑓(π‘₯) = π‘₯Β³ + 5π‘₯ βˆ’ 4 ; [ βˆ’3, βˆ’1 ]


4. β„Ž(π‘₯) = (π‘₯ βˆ’ 3)^1/3 + 4 ; [ 0, 2 ]


5. 𝑓(𝑑) = 3 cos 2𝑑 ; [ 1/6πœ‹,3/4πœ‹ ]


1
Expert's answer
2021-06-21T06:12:47-0400

Solution.

1.

"g(t)=2\sec{\frac{1}{2}t}, [-\frac{1}{3}π,\frac{1}{2}π]."

We find the critical point of g. To do so, take the derivative:

"g'(t)=\sec{\frac{t}{2}}\tan{\frac{t}{2}}."

Then "g'(t)=0" when"\sec{\frac{t}{2}}\tan{\frac{t}{2}}=0,\newline
\sec{\frac{t}{2}}=0\text{ or 
 }\tan{\frac{t}{2}}=0,\newline
t=2πk, k\in Z."

"t=0 \in [-\frac{1}{3}π,\frac{1}{2}π]."

Evaluate g at the critical points and on the endpoints of the interval:

"g(0)=2,\newline
g(-\frac{1}{3}π)=2\frac{1}{\cos{\frac{π}{6}}}=\frac{4}{\sqrt{3}},\newline
g(\frac{1}{2}π)=2\frac{1}{\cos{\frac{π}{4}}}=\frac{4}{\sqrt{2}}."

Therefore, g achieves its absolute minimum of 2 at t=0 and its absolute maximum of "\frac{4}{\sqrt{2}}" at "t=\frac{1}{2}π." .

2.

"g(x)=x^3+3x^2-9x, [-4,4]."

We find the critical point of g. To do so, take the derivative:

"g'(x)=3x^2+6x-9."

Then "g'(x)=0" when"3x^2+6x-9=0,\newline
x^2+2x-3=0,\newline
x=-3, \text{or } x=1."

Evaluate g at the critical points and on the endpoints of the interval:

"g(-3)=-27+27+27=27,\newline
g(1)=1+3-9=-5,\newline
g(-4)=-64+48+36=20,\newline
g(4)=64+48-36=76."

Therefore, g achieves its absolute minimum of -5 at x=1 and its absolute maximum of 76 at x=4.

3.

"f(x)=x^3+5x-4, [-3,-1]."

We find the critical point of f. To do so, take the derivative:

"f'(x)=3x^2+5."

Then "f'(x)=0" when"3x^2+5=0,\newline
x^2=-\frac{5}{3},\newline
x\in \varnothing."

Evaluate f only on the endpoints of the interval:

"f(-3)=-27+15-4=-16,\newline
f(-1)=-1+5-4=0.\newline"

Therefore, f achieves its absolute minimum of -16 at x=-3 and its absolute maximum of 0 at x=-1.

4.

"h(x)=(x-3)^{\frac{1}{3}}+4, [0,2]."

We find the critical point of h. To do so, take the derivative:

"h'(x)=\frac{1}{3}(x-3)^{-\frac{2}{3}}."

Then "h'(x)=0" when"\frac{1}{3}(x-3)^{-\frac{2}{3}}=0,\newline
x=3."

Evaluate h at the critical points and on the endpoints of the interval:

"h(3)=4,\newline
h(0)=-\sqrt[3]{3}+4,\newline
h(2)=-\sqrt[3]{1}+4=3.\newline"

Therefore, h achieves its absolute minimum of "-\sqrt[3]{3}+4" at x=0 and its absolute maximum of 4 at x=3.

5.

"f(t)=3\cos{2t}, [\frac{1}{6}π,\frac{3}{4}π]."

We find the critical point of f. To do so, take the derivative:

"f'(x)=-6\sin{2t}."

Then "f'(x)=0" when"-6\sin{2t}=0,\newline
\sin{2t}=0,\newline
t=\frac{πk}{2}, k\in Z."

"t=\frac{π}{2}\in [\frac{1}{6}π,\frac{3}{4}π]."

Evaluate f at the critical points and on the endpoints of the interval:

"f(\frac{π}{6})=3\cos{\frac{π}{3}}=1.5,\newline
f(\frac{3π}{4})=3\cos{\frac{3π}{2}}=0,\newline
f(\frac{π}{2})=3\cos{π}=-3.\newline"

Therefore, f achieves its absolute minimum of -3 at "t=\frac{π}{2}," and its absolute maximum of 1.5 at "t=\frac{π}{6}."